An elevator cab that weighs 27.0 kN moves upward. What is the tension in the cable if the cab's speed is (a) increasing at a rate of 1.28 m/s2 and (b) decreasing at a rate of 1.28 m/s2?
Answer #1

A.

when speed is increasing, then

using force balance

T = m*(g + a)

T = mg + ma

(m = W/g)

T = 27000 + (27000/9.81)*1.28 = 30.522 kN

B.

when speed is decreasing, then

using force balance

T = m*(g - a)

T = mg - ma

(m = W/g)

T = 27000 - (27000/9.81)*1.28 = 23.477 kN